OpenStudy (anonymous):

an equation of the tangent line to the function y^4=-5x^2-3x^2-3x+12. at the point (1,1). a)y=-7x+6 b)y=2x-3 c)y=-6x+7 d) -6x-8 e) y=-2x+1

myininaya (myininaya):

Have you found y' yet?

OpenStudy (anonymous):

i have but it is very messy. I was wondering if the y^4 might be a typo. i am not sure

myininaya (myininaya):

derivative of y^4 is 4y^3*y' by chain rule and power rule

myininaya (myininaya):

you could go ahead and combine the like terms -5x^2-3x^2 first before differentiating though

myininaya (myininaya):

that is if they were suppose to be like terms

OpenStudy (anonymous):

i am going to bet there is a typo

OpenStudy (anonymous):

i tried to do find the tangent line and i am not getting any of the choices.. i dont know

myininaya (myininaya):

well your problem has y^4=-5x^2-3x^2-3x+12 i'm kinda wondering if you meant y^4=-5x^3-3x^2-3x+12

OpenStudy (anonymous):

i copied it down in class... maybe i did...

myininaya (myininaya):

in that case your answer is listed

myininaya (myininaya):

let me see your derivative work for differentiating y^4=-5x^3-3x^2-3x+12

OpenStudy (anonymous):

the derivative of that is 4y^3=-15x^2-6x-3

myininaya (myininaya):

you must use chain rule for the y^4

myininaya (myininaya):

\[4y^3 \cdot y'=-15x^2-6x-3 \]

OpenStudy (anonymous):

oh so implicitly differentiate

myininaya (myininaya):

now you can replace the x's with 1 and the y's with 1 and solve for y'

myininaya (myininaya):

well whenever i see x and y together most always y is a function of x

myininaya (myininaya):

most always

myininaya (myininaya):

it could be the other way around are both x and y could be functions of time

myininaya (myininaya):

but yeah I assume here y is a function of x

OpenStudy (anonymous):

so is it c

myininaya (myininaya):

lol yes

OpenStudy (anonymous):

haha thanks!!!